# Hetero-EUCLID: Interpretable model discovery for heterogeneous hyperelastic materials using stress-unsupervised learning11footnotemark: 1

### v1

@@ 3.4 A comment on detecting noise
Table 11 provides statistical details of the 2D nodal residual force $f^{a}_{\text{res}}$ corresponding to each type of heterogeneity and noise-level considered in the study. Additionally, we provide statistical details of residual forces in a homogeneous specimen with the same thin square plate domain geometry in the same table for comparison. It is evident from the results that the dimensionless ratio: $\left(\dfrac{\mu_{f_{\mathrm{res}}}}{\sigma_{f_{\text{res}}}}\right)$ is a measure of noise in the displacement data, with a higher ratio indicating a higher displacement noise or a lack of heterogeneity in the domain. Another important observation is that alongside the previous ratio, the dimensionless ratio $\left(\dfrac{\sigma_{f_{\text{res}}}}{R_{\mathrm{max}}}\right)$ provides an indication of the existence of heterogeneity in the domain. If $\left(\dfrac{\sigma_{f_{\text{res}}}}{R_{\mathrm{max}}}\right)$$\leq 10^{-5}$ and $\left(\dfrac{\mu_{f_{\mathrm{res}}}}{\sigma_{f_{\text{res}}}}\right)$$\geq 1$ , it is reasonable to consider the domain to be nominally homogeneous. However, further increase in displacement noise can be expected to increase both ratios and make identifying heterogeneities (or the lack thereof) in the domain challenging. The statistical trends observed from $f^{a}_{\text{res}}$ are thus useful to characterize displacement noise and the existence of heterogeneities from actual experimental data.
@@ 4 Conclusion and outlook
As part of future work, we aim to extend Hetero-EUCLID to accommodate more complex history-dependent constitutive behaviors such as viscoelasticity to model a broader range of soft tissues and composites. As mentioned in Shi et al. [45], the paradigm of first segmenting the material and then identifying the constitutive parameters of each segment is highly dependent on the ability to have clean inter-material segments. However, this is not the case in situations involving functionally graded materials, and most practical experiments where the material segments have smooth (not sudden) transitions between properties. Furthermore, in the current study, we show that a high enough displacement noise would undermine the accuracy of the material segmentation, further reducing the accuracy of predicting each material’s properties. We therefore intend to explore frameworks that provide a position-dependent field of the material properties, as opposed to segment-specific properties. An experimental validation using DIC/ Digital Volume Correlation (DVC) data will also provide further insights into the merits, practical limitations, and improvements necessary in the proposed framework.
@@ A.1 Parsimony-driven Model Prior
where $\mathcal{N}_{+}$ denotes a non-negative truncated normal distribution with zero mean and variance $\nu_{s}\sigma^{2}$ (slab). The truncated normal, $\mathcal{N}_{+}$ is obtained from the standard normal distribution by restricting samples to non-negative values $(\theta_{i}\geq 0)$, a requirement that ensures physical admissibility. When $z_{i}=0$, the $i^{\text{th}}$ feature is considered inactive, meaning $\theta_{i}=0$, and its probability density reduces to a Dirac delta function (spike) centered at zero. Conversely, if $z_{i}=1$, the feature is active and $\theta_{i}$ is drawn from a truncated normal distribution $\mathcal{N}^{+}$ . Assuming that the elements of $\boldsymbol{\theta}$ are independent and identically distributed (i.i.d.), the joint prior for $\boldsymbol{\theta}$ can be expressed as
@@ A.2 Model Identification through Posterior Estimation
It is important to note that the direct analytical sampling from Equation 36 is challenging due to the presence of spike-slab priors, therefore Markov Chain Monte Carlo (MCMC) methods is deployed. In this work, we adopt a Gibbs sampling strategy [61] for posterior distribution estimation. The Gibbs sampler requires the conditional posterior distributions of all random variables. Closed-form expressions for these conditional posteriors were derived by Nayek et al. [54], and a concise summary of these results is provided in Table 2:
@@
Appendix B Optimal sub-sampling criterion at material interface boundaries

### v2

@@ 3.6 A comment on detecting noise
Specimen Type $R_{\mathrm{max}}$ [N] Noise Level $\lambda$ $\mu_{f_{\mathrm{res}}}$ [N] $\sigma_{f_{\mathrm{res}}}[N]$ $\dfrac{\mu_{f_{\mathrm{res}}}}{\mathit{\sigma_{f_{\text{res}}}}}$ $\dfrac{\mathit{\sigma_{f_{\text{res}}}}}{R_{\mathrm{max}}}$ CPU time (min) (Max. boundary (Tunable (Mean) (Standard force) scalar) deviation) Type 0 (Homogeneous) 904.56 Zero $2.38\times 10^{-5}$ $1.48\times 10^{-5}$ 1.608 $1.64\times 10^{-8}$ - Moderate 0.122 0.074 1.664 $8.13\times 10^{-5}$ - High 0.247 0.135 1.826 $1.49\times 10^{-4}$ - Type 1 (7(a)) 344.44 Zero 2.00 0.042 0.252 0.168 $7.31\times 10^{-4}$ 1.4 Moderate 2.00 0.085 0.247 0.343 $7.17\times 10^{-4}$ 2.0 High 1.50 0.134 0.243 0.552 $7.04\times 10^{-4}$ 2.5 Type 2 (7(b)) 879.28 Zero 1.50 0.065 0.380 0.172 $4.32\times 10^{-4}$ 11.3 Moderate 1.50 0.193 0.371 0.520 $4.22\times 10^{-4}$ 4.6 High 2.12 0.328 0.377 0.871 $4.28\times 10^{-4}$ 4.1 Type 3 (7(c)) 969.58 Zero 1.70 0.185 0.906 0.204 $9.34\times 10^{-4}$ 61.7 Moderate 1.70 0.338 0.902 0.375 $9.30\times 10^{-4}$ 77.6 High 1.80 0.507 0.888 0.571 $9.16\times 10^{-4}$ 113.2 Type 4 (7(d)) 562.26 Zero 1.50 0.095 0.472 0.202 $8.40\times 10^{-4}$ 108.5 Moderate 1.50 0.186 0.469 0.396 $8.34\times 10^{-4}$ 113.3 High 1.50 0.272 0.464 0.587 $8.25\times 10^{-4}$ 241.1 Type 5 (7(e)) 330.54 Zero 1.50 0.024 0.173 0.138 $5.24\times 10^{-4}$ 20.7 Moderate 1.50 0.067 0.171 0.390 $5.16\times 10^{-4}$ 32.0 High 1.80 0.114 0.171 0.666 $5.16\times 10^{-4}$ 39.1 Type 6 (7(f)) 795.96 Zero 2.00 0.095 0.642 0.149 $8.07\times 10^{-4}$ 3.6 Moderate 2.00 0.195 0.634 0.307 $7.97\times 10^{-4}$ 10.1 High 2.00 0.296 0.626 0.473 $7.87\times 10^{-4}$ 5.0
Table 9: Comparison of statistical measures (mean, standard deviation, and ratios) and computational time for each heterogeneous specimen type under different noise levels.
Table 9 provides statistical details of the 2D nodal residual force $f^{a}_{\text{res}}$ corresponding to each type of heterogeneity and noise-level considered in the study. Additionally, we provide statistical details of residual forces in a homogeneous specimen with the same thin square plate domain geometry in the same table for comparison. It is evident from the results that the dimensionless ratio: $\left(\mu_{f_{\mathrm{res}}}/\mathit{\sigma_{f_{\text{res}}}}\right)$ is a measure of noise in the displacement data, with a higher ratio indicating a higher displacement noise or a lack of heterogeneity in the domain. Another important observation is that alongside the previous ratio, the dimensionless ratio $\left(\mathit{\sigma_{f_{\text{res}}}}/R_{\mathrm{max}}\right)$ provides an indication of the existence of heterogeneity in the domain. If $\left(\mathit{\sigma_{f_{\text{res}}}}/R_{\mathrm{max}}\right)\leq 10^{-5}$ and $\left(\mu_{f_{\mathrm{res}}}/\mathit{\sigma_{f_{\text{res}}}}\right)\geq 1$, it is reasonable to consider the domain to be nominally homogeneous. However, a further increase in displacement noise can be expected to increase both ratios, making it challenging to identify heterogeneities (or the lack thereof) in the domain. The statistical trends observed from $f^{a}_{\text{res}}$ are thus useful to characterize displacement noise and the existence of heterogeneities from actual experimental data.
@@ 4 Conclusion and outlook
As part of future work, we aim to extend Hetero-EUCLID to accommodate more complex history-dependent constitutive behaviors, such as viscoelasticity, to model a broader range of soft tissues and composites, while accounting for spatial correlations in residual forces (see subsection 2.2). As mentioned in Shi et al. [2025], the paradigm of first segmenting the material and then identifying the constitutive parameters of each segment is highly dependent on the ability to have clean inter-material segments. However, clean inter-material segments are not available in situations involving functionally graded materials, and most practical experiments where the material segments have smooth (not sudden) transitions between properties. Furthermore, in the current study, we demonstrate that a sufficiently high displacement noise can undermine the accuracy of material segmentation, thereby further reducing the accuracy of predicting each material’s properties. We therefore intend to explore frameworks that provide a position-dependent field of the material properties, as opposed to segment-specific properties. An experimental validation using DIC/Digital Volume Correlation (DVC) data will also provide further insights into the merits, practical limitations, and necessary improvements in the proposed framework.
@@ A.1 Parsimony-driven Model Prior
where $\mathcal{N}_{+}$ denotes a non-negative truncated normal distribution with zero mean and variance $\nu_{s}\sigma^{2}$ (slab). The truncated normal ($\mathcal{N}_{+}$) is obtained from the standard normal distribution by restricting samples to non-negative values $(\theta_{i}\geq 0)$, a requirement that ensures physical admissibility. When $z_{i}=0$, the $i^{\text{th}}$ feature is considered inactive, meaning $\theta_{i}=0$, and its probability density reduces to a Dirac delta function (spike) centered at zero. Conversely, if $z_{i}=1$, the feature is active and $\theta_{i}$ is drawn from a truncated normal distribution $\mathcal{N}^{+}$ . Assuming that the elements of $\boldsymbol{\theta}$ are independent and identically distributed (i.i.d.), the joint prior for $\boldsymbol{\theta}$ can be expressed as
@@ A.2 Model Identification through Posterior Estimation
It is important to note that the direct analytical sampling from Equation 42 is challenging due to the presence of spike-slab priors; therefore, Markov Chain Monte Carlo (MCMC) methods are deployed. In this work, we adopt a Gibbs sampling strategy Casella and George [1992] for posterior distribution estimation. The Gibbs sampler requires the conditional posterior distributions of all random variables. Closed-form expressions for these conditional posteriors were derived by Nayek et al. [2021], and a concise summary of these results is provided in Table 10:
@@
Appendix B Moran’s I test for assessing spatial correlations in residual force
We conduct a series of spatial correlation studies separately on the $x$, $y$, and $z$ degrees of freedom of the residual force vector $\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}=\boldsymbol{A\theta-b}$, with $\boldsymbol{A}$, $\boldsymbol{\theta}$ and $\boldsymbol{b}$ as defined in Equation 22 of the main paper. Further, Figure 5 of the main paper illustrates the process to obtain $\left(\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}\right)_{x}$, $\left(\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}\right)_{y}$ and $\left(\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}\right)_{z}$, which are the $x$, $y$ and $z$ components of $\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}$ vector. To examine spatial correlations in different components of the residual force vector, we make use of Moran’s index ($I$).
Given any residual vector $\boldsymbol{\phi}$, the Moran’s index ($I$) [Moran, 1950] is defined as follows:
$I=\frac{N}{W}\left[\frac{\sum_{i=1}^{N}\sum_{j=1}^{N}(\phi_{i}-\mu)W_{ij}(\phi_{j}-\mu)}{\sum_{i=1}^{N}(\phi_{i}-\mu)^{2}}\right]$ (45)
where $\phi_{i}$ denotes the $i^{\text{th}}$ component of the residual vector $\boldsymbol{\phi}$, $\mu=(1/N)\sum_{i=1}^{N}\phi_{i}$ and $\boldsymbol{W}$ is a weight matrix indicating connectivity between nodes, i.e., $W_{ij}$ defines the connection between the $i^{\text{th}}$ and $j^{\text{th}}$ component of the residual vector and $W_{ii}=0$ (no summation), i.e, the weight matrix has all diagonal entries set to zero.
In this study, $\boldsymbol{\phi}=\left(\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}\right)_{x}$, $\left(\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}\right)_{y}$ or $\left(\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}\right)_{z}$ (various components of the residual force vector), and $N$ would be equal to the number of nodes in the finite element mesh. $\boldsymbol{W}$ is constructed using the finite element connectivity matrix by setting $W_{ij}=1$, only if the $i^{\text{th}}$ and $j^{\text{th}}$ nodes are connected, and $W_{ij}=0$ otherwise. For a residual vector ($\boldsymbol{\phi}$) that is sampled from an independent and identical distribution (i.i.d.), there would be no correlations between different components of the residual vector. For such cases, Moran’s index $I$ would be close to zero. Conversely, a Moran’s index close to 0 indicates the lack of spatial correlations in the residual vector, and a higher magnitude of the Moran’s index indicates stronger spatial correlations between different components of the residual vector.
B.1 Test study 1 - Evaluating I for residual force vector sampled from an i.i.d.
As the first benchmark study for Moran’s index, a synthetic residual force vector is sampled from an independent and identically distributed (i.i.d.) normal distribution (Equation 46).
$\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}=\boldsymbol{A\theta-b}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ (46)
Then, $I$ is evaluated for the residual vectors $\left(\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}\right)_{x}$, $\left(\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}\right)_{y}$, and $\left(\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}\right)_{z}$ using Equation 45. As shown in Figure 17, the resulting Moran’s index values remain close to zero for all components of the residual force vector. This is consistent with sampling the residual force vector from an i.i.d.
Figure 17: Residual force samples for the $x$, $y$, and $z$ degrees of freedom sampled from an i.i.d. $\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ distribution for Moran’s index (I) evaluation, illustrating the expected near-zero Moran’s index for uncorrelated data.
B.2 Test study 2 - Evaluating I for residual force vector sampled from a correlated normal distribution.
As the next benchmark study for Moran’s index, we generate a residual force vector from a spatially correlated zero-mean Gaussian distribution:
$\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}=\boldsymbol{A\theta-b}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{\Sigma})$ (47)
where $\boldsymbol{\Sigma}$ denotes the covariance matrix. The entries of $\boldsymbol{\Sigma}$ are defined through an exponential kernel so that nodes that are closer in space share more strongly correlated noise.
$\Sigma_{pq}=\exp\left(-\frac{d_{pq}}{\ell_{c}}\right),\qquad d_{pq}=\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{2}\quad\text{where},\quad p,q\in\{1,\dots,n_{n}\}$ (48)
where $d_{pq}$ represents the Euclidean distance between nodes $p$ and $q$, and $\ell_{c}$ denotes the correlation length. For the $50~\text{mm}\times 50~\text{mm}\times 1~\text{mm}$ hyperelastic specimen considered here, we choose $\ell_{c}=1.0~\text{mm}$ equivalent to 2% of the specimen size of in–plane dimension.
Figure 18: Residual force samples for the $x$, $y$, and $z$ degrees of freedom drawn from a cross-correlated Gaussian distribution $\mathcal{N}(\boldsymbol{0},\boldsymbol{\Sigma})$, illustrating the increase in Moran’s index (I) due to spatial correlation embedded in the covariance structure.
The resulting $I$ values for $\left(\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}\right)_{x}$, $\left(\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}\right)_{y}$, and $\left(\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}\right)_{z}$ (Figure 18) are noticeably higher than those observed in test study 1 (B.1), although the magnitudes of the residual forces are similar in both test studies. Thus, from test studies 1 and 2, we can confirm that the $I$ values for the residual forces increase with an increase in spatial correlation. Additionally, we obtain benchmarks to ascertain the extent of spatial correlation for a given residual force vector, i.e., $I$ values closer to those of test study 1 indicate the absence of spatial correlations. In contrast, $I$ values closer to those of test study 2 indicate the presence of significant spatial correlations.
B.3 Test study 3 - Effect of uncorrelated displacement noise
In this final test study, we calculate $I$ values for residual forces obtained from simulated biaxial tension tests on a thin, square-shaped, homogeneous hyperelastic specimen. The finite element simulation discretizes the geometry using 10,158 finite elements (C3D6 type). The hyperelastic material was prescribed a NH2 constitutive model with strain-energy given by $W_{\mathrm{NH2}}=\mu(\tilde{I}_{1}-3)+K(J-1)^{2}$. The residual force vector $\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}=\boldsymbol{A\theta-b}$ is obtained using the ground-truth material model $\boldsymbol{\theta}$ that was used in the simulations ($\mu=$ 5.4 and $K=$ 15.0).
Figure 19: Moran’s index ($I$) score for a homogeneous hyperelastic material subject to biaxial tension, using 10,158 elements under an uncorrelated noise ($\sigma_{u}$) with three noise levels as ($2\times 10^{-3}\>,\>5\times 10^{-3}\>\text{and}\>1\times 10^{-2}$) to the displacement field.
An uncorrelated displacement noise $\sigma_{u}$ is added to the displacements obtained from these simulations. The addition of uncorrelated displacement noise ($\sigma_{u}=2\times 10^{-3},~5\times 10^{-3},~1\times 10^{-2}$) will only affect the $\boldsymbol{A}$ matrix, which is computed using the nodal displacement data and the connectivity table. The vector $\boldsymbol{b}$ is set to $\boldsymbol{0}$. The resulting $I$ values for $\left(\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}\right)_{x}$, $\left(\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}\right)_{y}$, and $\left(\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}\right)_{z}$ (see Figure 5) are shown in Figure 19. The $I$ values, being an order of magnitude higher than those encountered in test study 1 (B.1), indicate similar significant spatial correlations in residual forces for all noise cases.
Thus, the spatial correlations in the residual forces observed in test study 3 (B.3) indicate that it would be inaccurate to sample the force vector $\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}$ from an independent and identical distribution (i.i.d.) as was done in the likelihood equation (Equation 33 in the Bayesian-EUCLID paper [Joshi et al., 2022] and Equation 24 in this paper).
Appendix C Sampling residual forces from an i.i.d. Gaussian distribution is equivalent to solving OLS and minimizing MSE.
Given the non-negligible spatial correlations between components of the residual force vector ($\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}$), in B we have established the inaccuracy of Equation 33 in the Bayesian-EUCLID paper [Joshi et al., 2022] and Equation 24 in the current Hetero-EUCLID paper. However, in this section, we show that many conventional inverse material identification approaches implicitly make the same incorrect assumption. Specifically, all data-driven approaches that employ the Ordinary Least Squares (OLS) (such as the equilibrium-gap methods [Flaschel et al., 2021b, Haustrate et al., 2024]) or that minimize the Mean-Squared Error (MSE) of residual forces (such as NN-based stress-unsupervised frameworks [Thakolkaran et al., 2022, Lourenço et al., 2024, Meng et al., 2025, Shi et al., 2025]) implicitly assume a spatially uncorrelated residual force vector. To prove this, we first establish the equivalences between obtaining the material parameters $\boldsymbol{\theta}$ by maximizing the i.i.d. based likelihood (Equation 24 of the Hetero-EUCLID framework), and obtaining $\boldsymbol{\theta}$ using OLS or by minimizing an MSE-based objective function.
In the Bayesian-EUCLID [Joshi et al., 2022] framework to model hyperelastic materials, we have the residual forces given by a linear function of the material parameters ($\boldsymbol{\theta}$), i.e, $\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}=\boldsymbol{A\theta-b}$. However, in general, $\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}$ can also be a non-linear function of the material parameters ($\boldsymbol{\theta}$), as would happen in the case of using neural networks to obtain the residual force [Thakolkaran et al., 2022, Shi et al., 2025] or when dealing with non-hyperelastic materials [Flaschel et al., 2022, Lourenço et al., 2024]. In the Bayesian-EUCLID formulation, if we were to impose a more generic likelihood (accounting for spatial correlations in the residual force) for satisfying the weak form of the linear momentum balance, we would have:
$\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}\sim\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma}),$ (49)
where $\boldsymbol{\mu}$ is the mean vector and $\boldsymbol{\Sigma}$ is the covariance matrix.
Equation 49 can be written as:
$P(\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}})$ $=\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$
$=K\exp\left(-\frac{1}{2}[\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}-\boldsymbol{\mu}]^{T}\boldsymbol{\Sigma}^{-1}[\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}-\boldsymbol{\mu}]\right)$ (50)
where $K$ is a proportionality constant that collects all factors independent of $\boldsymbol{\theta}$.
To discover the material parameters ($\boldsymbol{\theta}^{*}$), we would have to maximize the likelihood of satisfying the weak form of linear momentum balance Equation 50. That is,
$\boldsymbol{\theta}^{*}=\operatorname*{argmax}_{\theta}\>P(\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}})$ (51)
Maximizing the likelihood (Equation 51) is the same as maximizing its logarithm:
$\boldsymbol{\theta}^{*}=\operatorname*{argmax}_{\theta}\>P(\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}})=\operatorname*{argmax}_{\theta}\>\log P(\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}})$ (52)
The logarithm of the likelihood can be further simplified as:
$\log P(\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}})=\log K-\frac{1}{2}[\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}-\boldsymbol{\mu}]^{T}\boldsymbol{\Sigma}^{-1}[\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}-\boldsymbol{\mu}]$ (53)
Since $K$ is independent of $\boldsymbol{\theta}$, Equation 52 can be written as
$\boldsymbol{\theta}^{*}=\operatorname*{argmin}_{\theta}\>[\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}-\boldsymbol{\mu}]^{T}\boldsymbol{\Sigma}^{-1}[\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}-\boldsymbol{\mu}]$ (54)
In Equation 49, if we were to set $\boldsymbol{\mu}=\boldsymbol{0}$ (zero-mean) and $\boldsymbol{\Sigma}=\sigma^{2}\boldsymbol{I}$ (i.i.d.), we would recover the likelihood for satisfying the weak form of linear momentum balance used in the Bayesian-EUCLID paper [Joshi et al., 2022]. For $\boldsymbol{\mu}=\boldsymbol{0}$ and $\boldsymbol{\Sigma}=\sigma^{2}\boldsymbol{I}$, Equation 54 will simplify to:
$\boldsymbol{\theta}^{*}=\operatorname*{argmin}_{\theta}\>\left\lVert\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}\right\rVert_{2}^{2}$ (55)
Equation 55 describes the process of minimizing the Mean Squared Error (MSE) of the residual forces ($\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}$), and inherently assumes that the expected value of the residual forces is zero, i.e, $\boldsymbol{\mu}=\boldsymbol{0}$, and that there are no spatial correlations in the residual forces, i.e, $\boldsymbol{\Sigma}=\sigma^{2}\boldsymbol{I}$. Many works that use neural networks to surrogate the constitutive model of materials in a stress-unsupervised manner use the MSE loss of residual forces to train their neural networks [Thakolkaran et al., 2022, Shi et al., 2025, Meng et al., 2025]. Thus, these works also implicitly assume that the residual forces ($\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}$) are expected to be $\boldsymbol{0}$ and have no spatial correlations. In order to correctly account for spatial correlations in the residual force, these works ([Thakolkaran et al., 2022, Shi et al., 2025, Meng et al., 2025]) should have used Equation 54 as their loss function, instead of Equation 55.
Further, in Equation 55, if we substitute $\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}=\boldsymbol{A\theta-b}$, we obtain:
$\boldsymbol{\theta}^{*}=\operatorname*{argmin}_{\theta}\>\left\lVert\boldsymbol{A\theta-b}\right\rVert_{2}^{2}$ (56)
Equation 56 defines the Ordinary Least Squares (OLS) solution to obtain the material parameters $\boldsymbol{\theta}^{*}$, and inherently assumes that the expected value of the residual forces is zero, i.e, $\boldsymbol{\mu=0}$ and that there are no spatial correlations in the residual forces, i.e, $\boldsymbol{\Sigma}=\sigma^{2}\boldsymbol{I}$. This approach is commonly used to obtain material parameters in equilibrium gap-based methods ([Haustrate et al., 2024, Flaschel et al., 2021b]), which by extension, also implicitly assume lack of spatial correlations in the residual forces ($\boldsymbol{f}_{\mathrm{res,}\boldsymbol{\theta}}$).
Appendix D Optimal sub-sampling criterion at material interface boundaries